1.Utopias residents are endowed with 100 labor units and 75 capital units. Idealands residents are endowed with 75 labor units and 50 capital units. which nation is relatively labor abundant and which is relatively capital abundant?
2. Suppose that, at the current level of production, sandwiches require 1 labor unit and 0.5 capital units to produce, while bicycles require 0.9 labor units and 0.9 capital units. Based on this information, which of the two goods is relatively capital intensive in its production process, and which is relatively labor intensive?
3.Answer the following using your answers to question 1 and 2. Suppose utopias and idealands residents begin to trade. Which good will utopias residents export, and which good will utopias residents import? explain your answer in the context of the Heckscher-ohlin theorem.
5.Based on your answer to question 3, explain how free trade between the residents of utopia and idealand benefits or harms workers and the capital holders in the two nations.
1
Expert's answer
2017-02-11T08:32:13-0500
1. The calculations of the ratio of capital to labor Utopia are 100/75 = 1.33, respectively Idealand - 75/50 = 1,5. From this we can conclude that Idealand is relatively labor abundant and Utopia is relatively capital abundant.
2. Display this mathematically as sandwich company has a production model that is twice as labor intensive as capital (0.5 / 1 = 0.5) in the preparation of the bike evenly labor and capital intensive (0.9 / 0.9 = 1). So bikes are relatively capital-intensive in their production process while sandwiches are relatively labor-intensive in their production process.
3. Find output quantities, assuming full employment of both factors in both countries:
Utopia: 0.5 S + 0.9 B = 75, 1 S+ 0.9 B = 100 S = 50, B = 56
Idealand: 0.5 S + 0.9 B = 50, 1 S+ 0.9 B = 75 S = 28, B = 40
Based on the fact that Utopia is relatively capital abundant and bicycles are relatively capital intensive in their production process, Utopia has a comparative advantage in bicycles and will export them to Idealand.
Based on the fact that Idealand is relatively labor abundant and sandwiches are relatively labor intensive in their production process, Utopia has a comparative advantage in bicycles and will export them to Idealand.
5. With free trade will rise real returns to the holders of the relatively abundant factor in Utopia, capital, while returns to the holders of the relatively scarce factor, labor, will decline. in Idealand occurs the opposite.